# Privacy-Preserving Federated Learning Scheme with Mitigating Model Poisoning Attacks: Vulnerabilities and Countermeasures

### v1

@@ 6.1 Security Analysis
6.2 Performance Comparison
Our scheme has lower computational and communication burdens compared with ShieldFL. We compare the complexity of the two schemes in their split processes and the comparison results are shown in Table III and Table IV.
TABLE V: Computational costs of the encryption algorithm in ShieldFL and our scheme.
Schemes Encryption Decryption Homo. Addition Homo. Multiplication
ShieldFL $penc$ $pdec$ $padd$ $pmul$
$O(N(mN\cdot mul+add))$ $O(N((N+1)\cdot mul+add))$ $O(N^{2}\cdot mul)$ $O(2N\cdot rnd+(4mN^{2}+9N^{2}+6N)mul+10N\cdot add)$
Our scheme $fenc$ $fdec$ $fadd$ $fmul$
$O(n_{f}q\cdot(2mul+3add))$ $O(n_{f}q\cdot(mul+3add))$ $O(n_{f}q\cdot add)$ $O(n_{f}q\cdot(5mul+2add))$
TABLE VI: Communication costs of the encryption algorithm in ShieldFL and our scheme
Schemes Homo. multiplication Ciphertext size
ShieldFL $C_{pmul}:O(5N^{2})$ $C_{p}:O(N^{2})$
Our scheme 00 $C_{f}:O(p)$
  • In Setup process of both schemes, KC and $S_{1}$ distribute keys and initial model, respectively. The computation complexities of ShieldFL and our scheme are $O(\varepsilon+(n+1)\cdot ks+l\cdot penc+(2n+1)l\cdot pdec)$ and $O(\varepsilon^{\prime}+(n+1)n_{f}\cdot ks+\tau\cdot fenc+(2n+1)\tau\cdot fdec)$, respectively, where $ks$, and $penc/fdec$, and $pdec/fdec$ represent the running time when implementing $\mathbf{PHE.KeySplit},$ $\mathbf{PHE.Enc}/$$\mathbf{FHE.Enc},$ and $\mathbf{PHE.Dec}/$$\mathbf{FHE.Dec}$, respectively. The communication complexities for ShieldFL and our scheme are $O((2n+2)\varepsilon+2nlC_{p})$ and $O((2n+2)n_{f}\varepsilon^{\prime}+2n\tau C_{f})$, respectively, where $C_{p}$/$C_{f}$ represents the ciphertext sizes of the two-trapdoor PHE/FHE.
  • In Local Training process of both schemes, $n$ users train their models locally, encrypt their local gradients, and transmit them to the server. The computation complexities of ShieldFL and our scheme are $O(nlT\cdot train+nlT\cdot penc))$ and $O(nlT\cdot train+n\tau T\cdot fenc))$, respectively, where $T$ is the total training iterations. The communication complexities of ShieldFL and our scheme are $O(nlTC_{p})$ and $O(n\tau TC_{f})$, respectively.
  • In Privacy-Preserving Defense process, the computation and communication costs of both schemes mainly come from the protocols $\mathbf{SecJudge}/\mathbf{ESecJudge}$ and $\mathbf{SecCos}/\mathbf{ESecCos}$. For ShieldFL, the computation and communication complexities for executing the two protocols are $O(4nlT(N\cdot rnd+penc+9pdec+padd+N^{2}\cdot pmul))$ and $O(nT(10l+9)C_{p})$, respectively, where $rnd,padd$ and $pmul$ represent the running time of generating a random noise, a homomorphic addition, and a homomorphic multiplication, respectively. For the enhanced scheme, the computation and communication complexities for executing the two protocols are $O(n\tau T(3q\cdot rnd+3fenc+9fdec+5fadd+3fmul))$ and $O(nT(2\tau+4)C_{f})$, respectively, where $fadd/fmul$ represent the running time of a fully homomorphic addition/multiplication.
  • In Model Update process, all users interact with $S_{1}$ to obtain the decrypted aggregated gradient and then update their local models. For ShieldFL, the computation and communication complexities of this process are $O(nlT(3\cdot pdec+update))$ and $O(2nlTC_{p})$, respectively. For our scheme, the computation and communication complexities are $O(3n\tau T\cdot fdec+nlT\cdot update)$ and $O(2n\tau TC_{f})$, respectively.
The total computational costs of ShieldFL and our scheme are $O(\varepsilon+(n+1)\cdot ks+nlT(8penc+36pdec+4padd+4N^{2}\cdot pmul+4N\cdot rnd+train+update)$ and $O(\varepsilon^{\prime}+(n+1)n_{f}\cdot ks+n\tau T(7fenc+9fdec+5fadd+3fmul+3q\cdot rnd)+nlT\cdot(train+update))$, respectively. Table V presents the computational costs of $penc,pdec,padd,pmul,fenc,fdec,fadd,fmul$ related to general addition $add$ and general multiplication $mul$ operations. Thus, the total computational costs of ShieldFL and our scheme are $O(\varepsilon+(n+1)\cdot ks+nlT(36N^{4}\cdot mul+40N^{3}\cdot add+8N^{3}\cdot rnd+train)$ and $O(\varepsilon^{\prime}+(n+1)n_{f}\cdot ks+n\tau T(35n_{f}q\cdot mul+59n_{f}q\cdot add+9q\cdot rnd)+nlT\cdot train)$, respectively. The total communication costs of ShieldFL and the enhanced scheme are $O(2n\varepsilon+13nlTC_{p})$ and $O(2nn_{f}\varepsilon^{\prime}+5n\tau TC_{f})$, respectively. Table VI presents the communication costs of homomorphic multiplication operations and the ciphertext sizes of the two schemes. Our scheme’s homomorphic multiplication operation is performed locally, and thus its communication cost is 00. The homomorphic multiplication protocol in ShieldFL requires interactions between the two servers, during which five ciphertexts are transformed. Thus, its communication cost is $O(5N^{2})$. Therefore, the total communication costs of ShieldFL and the enhanced scheme are $O(2n\varepsilon+13nlTN^{2})$ and $O(2nn_{f}\varepsilon^{\prime}+5n\tau Tq)$, respectively.

### v2

@@ 6.1 Security Analysis
6.2 Convergence Analysis
We provide a detailed theoretical convergence analysis of the proposed PBFL scheme in Appendix C. Specifically, we decompose the aggregation error into three components—(i) residual Byzantine bias, (ii) stochastic gradient noise, and (iii) gradient diversity—arising respectively from malicious updates, local sampling noise, and data heterogeneity. We then formally show that, under standard smoothness and bounded-variance assumptions, the proposed algorithm achieves the same sub-linear convergence rate $\mathcal{O}(1/\sqrt{T})$ as classical FedAvg, up to an additive error term that is proportional to the Byzantine fraction $\beta$ and decreases as the number of trusted participants increases. This result theoretically confirms the robustness and efficiency of our defense mechanism.
6.3 Performance Analysis
We provide a detailed performance comparison of computational and communication complexities between our proposed scheme and ShieldFL, which is presented in Appendix D. The analysis covers all major phases including Setup, Local Training, Privacy-Preserving Defense, and Model Update. Our results show that, due to ciphertext computation simplicity and cryptographic primitive efficiency, our scheme significantly outperforms ShieldFL in both computation and communication efficiency. Specifically:
  • Computational Complexity: Our scheme reduces asymptotic encryption, decryption, and homomorphic operation costs by a factor related to the smaller ciphertext dimension ($n_{f}q\ll N^{2}$). This translates to a lower overhead in all training rounds.
  • Communication Cost: Since homomorphic multiplication in our secure computation methods $\mathbf{ESecCos}$ and $\mathbf{ESecJudge}$ is executed locally without inter-party interaction, its communication cost is effectively zero. In contrast, ShieldFL incurs significant overhead due to interactive homomorphic multiplication in their secure computation protocols.
6.4 Discussion and Practical Considerations
While our enhanced scheme achieves significant improvements in privacy and efficiency, several practical considerations remain for real-world FL deployments. This section discusses two key aspects: the feasibility of decentralizing the Key Center and the challenges of defending against adaptive and malicious adversaries.
(1) Decentralizing the Key Center: Our current design assumes a trusted centralized Key Center for key management and distribution. This assumption, however, may not always hold in heterogeneous or cross-silo FL environments where participants belong to different administrative domains with varying trust levels. To mitigate this limitation, decentralized approaches such as threshold cryptography and Distributed Key Generation (DKG) protocols can be adopted. These techniques distribute key management responsibilities among multiple independent parties, thereby eliminating the single point of trust and reducing the risk of key compromise. In addition, blockchain-based solutions and secure multiparty computation (MPC) protocols could be integrated to provide transparent, auditable, and tamper-resistant key management. While these decentralized mechanisms can enhance robustness and trustworthiness, they inevitably introduce challenges such as increased communication overhead and coordination complexity, which require further study.
(2) Defending Against Adaptive and Malicious Threats: Our security analysis primarily considers malicious clients and at most one corrupted (malicious) server, ensuring the privacy of honest clients’ data under the assumption that the two servers do not collude. However, in real-world FL deployments, adversaries may behave both maliciously and adaptively, dynamically deviating from the protocol to launch sophisticated attacks such as adaptive model poisoning or adaptive inference.
We explicitly evaluate our framework under an adaptive model poisoning threat, where the adversary dynamically adjusts its malicious gradients based on the gradients of benign clients before injecting them into the training process. This adaptive strategy makes detection substantially more difficult. Our experimental results demonstrate that the proposed framework can effectively mitigate this class of adaptive attacks. Beyond model poisoning, more advanced adaptive threats have been reported, including adaptive backdoor attacks that iteratively modify backdoor triggers to evade defenses and adaptive inference attacks that progressively extract private information from model updates. These attacks evolve over training rounds and present challenges that our current design does not fully address.
Future work will focus on strengthening the framework against these stronger adversarial models by extending robustness beyond poisoning to cover adaptive backdoor and inference-based attacks. Potential directions include developing adaptive anomaly detection frameworks that dynamically adjust detection thresholds or defense strategies based on adversarial behaviors observed during training (e.g., online learning-based detectors that evolve as attacks change) and incorporating temporal consistency checks and cross-round validation mechanisms to detect adversaries that gradually adapt their malicious updates to evade static defenses. Furthermore, leveraging ensemble or diversity-based defenses (where multiple defense mechanisms operate in parallel and dynamically reconfigure based on the detected threat level) could make it substantially harder for adaptive attackers to circumvent all defenses simultaneously. Beyond adaptive attacks, we also plan to address the threat of malicious servers that compromise computation correctness by integrating verifiable computation or zero-knowledge proofs to guarantee both privacy and correctness even when a server is malicious.